OpenStudy (anonymous):

Solve the following radical equation: equation Select one: a. Take the square root of each side to solve for x2 b. Square each side of the equation to remove the radicand c. Subtract 7 from each side d. Subtract 1 from each side

OpenStudy (anonymous):

\[\sqrt{x-1}=x-7\]

OpenStudy (anonymous):

square both sides (carefully) to get \[x-1=x^2-14x+49\] then solve the quadratic equation

OpenStudy (anonymous):

oh i see the question only asks what to do first go with b. Square each side of the equation to remove the radicand